OpenStudy (anonymous):

the formula for converting degrees celsius (c) to degrees fahrenheit (f) is f= 9/5c + 32 . solve the formula for degrees celsius (c) . then find the temperature in degrees celsius (c) when the temperature in degrees fahrenheit (f) is 14. (please show work so i can understand)

OpenStudy (anonymous):

f=9/5c+32 f-32=9/5c 5/9(f-32)=c